OpenStudy (anonymous):

A square pyramid has a base area of 36 ft2. The pyramid’s total surface area is 144 ft2. What is the area of each triangular face? A. 108 ft2 B. 54 ft2 C. 36 ft2 D. 27 ft2

OpenStudy (anonymous):

plz help

OpenStudy (solomonzelman):

is the pyramid's bottom base an equilateral triangle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think

OpenStudy (solomonzelman):

Lets the (surface area of each) side=x so 36+3x=144 see how I get this equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's not that though

OpenStudy (anonymous):

look at the answers

OpenStudy (solomonzelman):

Cool, your smart, solve for x to find each side

OpenStudy (solomonzelman):

@angel12310, solve for x, you will get one of the choices.

OpenStudy (anonymous):

what i don't know how i'm not good at this only fractions

OpenStudy (solomonzelman):

36+3x=144 (36 is base, and 3x is all three side-faces of the pyramid.) subtract 36 from both sides and divide both sides by 3.

OpenStudy (anonymous):

how???